# thanks: Talento-CM fellow

### v1

@@ IV.7 Comparison with other cluster surveys
  • Probe $N$ provides a weak constraint on $M_{\nu}$, $\sigma^{N}(M_{\nu})=5.458\ (1.095)$ eV in the self-calibration (known SR) scenario. The constraint is mainly contributed by $P$, with $\sigma^{P}(M_{\nu})=0.293\ (0.223)$ eV. However, when $N$ is added, the constraint can be improved by a factor of 2.31 (6.56), resulting in $\sigma^{N+P}(M_{\nu})=$ 0.127 (0.034) eV.
@@
  • Having knowledge of the SR parameters has a significant impact on $\sigma^{N}(M_{\nu})$ and consequently affects $\sigma^{N+P}(M_{\nu})$. While partial knowledge leads to a small improvement in $\sigma^{N+P}(M_{\nu})$ from $0.127$ eV to $0.125$ eV, having perfect knowledge results in a substantial improvement to $0.034$ eV. This corresponds to a $2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
@@
  • Increasing $z_{max}$ from 1.5 to 2.0 will detect more clusters, increasing the count from $4\times 10^{5}$ to $4.3\times 10^{5}$. While this change has a very small impact on $\sigma^{P}(M_{\nu})$, improving it by a factor of 1.02 (1.02) in the self-calibration (known SR) scenario, it significantly strengthens $\sigma^{N}(M_{\nu})$ by a factor of 1.64 (1.20). As a result, it also strengthens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.08 (1.03), leading to $\sigma^{N+P}(M_{\nu})=0.118(0.033)$ eV.
@@
  • Among the uncertainties we considered, the redshift uncertainty $\sigma_{z}$ has the largest impact on $\sigma^{P}(M_{\nu})$, relaxing the constraint by a factor of 2.23 (2.13) when increasing $\sigma_{z}$ from 0.001 to 0.01. Consequently, this leads to a relaxation in $\sigma^{N+P}(M_{\nu})$ by a factor of 1.90 (2.15) in the self-calibration (known SR) scenario.
@@
  • We also examine the impact of uncertainties in the HMF and bias. The HMF uncertainties primarily affect $\sigma^{N}(M_{\nu})$, while the bias uncertainties mainly affect $\sigma^{P}(M_{\nu})$. However, the effects of both uncertainties on $\sigma^{N+P}(M_{\nu})$ are comparable, leading to a relaxation of the constraints by a factor of approximately 2 (4) in the self-calibration (known SR) scenario
@@
  • Finally, we investigate the impact of the GISDB effect, which has been omitted in the previous work. This effect compensates for the suppression of the power spectrum caused by neutrinos on small scales, thereby relaxing the constraint on neutrinos. However, taking into account the degeneracy between $M_{\nu}$ and bias, as well as the addition of the probe $N$, our results demonstrate that considering this effect actually tightens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.50 (2.21) in the self-calibration (known SR) scenario.
@@
In summary, CSST clusters are expected to provide a constraint on $M_{\nu}$ of $0.127$ eV. With perfect knowledge of the SR parameters, this constraint can be strengthened to $0.034$ eV, enabling a $2\sigma$ detection of the minimum neutrino mass of 0.06 eV. Recently, Lin et al. [86] presented a forecast using mock data from the CSST photometric galaxy clustering and cosmic shear surveys, which yielded an upper limit for $M_{\nu}$ of approximately $0.23$ eV at a 68% C.L. By combining clusters and galaxies, it is expected that CSST will significantly strengthen the constraint on the total neutrino mass. Moreover, the emergence of new probes such as voids [87] and Minkowski functions [88, 89] has opened up additional avenues for studying neutrino masses. By combining these innovative probes with traditional methods, it is anticipated that in the near future it will be possible to accurately constrain neutrino masses and distinguish between different mass hierarchies.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China Grants No. 12173036, 11773024, 11653002, 11421303, by the National Key R&\&&D Program of China Grant No. 2021YFC2203100, by the China Manned Space Project Grant No. CMS-CSST-2021-B01, by the Fundamental Research Funds for Central Universities Grants No. WK3440000004 and WK3440000005, and by the CAS Interdisciplinary Innovation Team.

### v2

@@ IV.7 Comparison with other forecasts
  • $N$ provides a weak constraint on $M_{\nu}$, $\sigma^{N}(M_{\nu})=5.458\ (1.095)$ eV in the self-calibration (known SR) scenario. The constraint is dominated by $P$, with $\sigma^{P}(M_{\nu})=0.293\ (0.223)$ eV. However, when $N$ is added, the constraint can be improved by a factor of 2.31 (6.56), resulting in $\sigma^{N+P}(M_{\nu})=$ 0.127 (0.034) eV.
@@
  • Knowledge of the SR parameters has a significant impact on $\sigma^{N}(M_{\nu})$ and consequently affects $\sigma^{N+P}(M_{\nu})$. While partial knowledge leads to a small improvement in $\sigma^{N+P}(M_{\nu})$ from $0.127$ eV to $0.125$ eV, having perfect knowledge results in a substantial improvement to $0.034$ eV. This corresponds to a $\sim 2\sigma$ detection of the minimum neutrino mass of $0.06$ eV.
@@
  • Increasing $z_{max}$ from 1.5 to 2.0 will result in more detected clusters, increasing the total count from $4\times 10^{5}$ to $4.3\times 10^{5}$. While this change has a very small impact on $\sigma^{P}(M_{\nu})$, improving it by a factor of 1.02 (1.02) in the self-calibration (known SR) scenario, it does strengthen $\sigma^{N}(M_{\nu})$ by a factor of 1.64 (1.20). As a result, it strengthens $\sigma^{N+P}(M_{\nu})$ a little bit, by a factor of 1.08 (1.03), leading to $\sigma^{N+P}(M_{\nu})=0.118(0.033)$ eV.
@@
  • Among the systematics we considered, the redshift uncertainty $\sigma_{z}$ has the largest impact on $\sigma^{P}(M_{\nu})$, relaxing the constraint by a factor of 2.23 (2.13) when increasing $\sigma_{z}$ from 0.001 to 0.01. Consequently, this leads to a degradation in $\sigma^{N+P}(M_{\nu})$ by a factor of 1.90 (2.15) in the self-calibration (known SR) scenario.
@@
  • We also examine the impact of uncertainties in the HMF and cluster bias. The HMF uncertainties primarily affect $\sigma^{N}(M_{\nu})$, while the bias uncertainties only affect $\sigma^{P}(M_{\nu})$. However, the effects of both uncertainties on $\sigma^{N+P}(M_{\nu})$ are comparable, leading to a degradation of the constraints by a factor of approximately 2 (4) in the self-calibration (known SR) scenario
@@
  • Finally, we investigate the impact of the GISDB, which has been missed in previous studies. This effect compensates for the suppression of the power spectrum caused by massive neutrinos on small scales, thereby one would expect it relaxes the constraint on neutrino mass. However, taking into account the degeneracy between $M_{\nu}$ and bias, as well as the addition of the probe $N$, our results show that considering this effect actually tightens $\sigma^{N+P}(M_{\nu})$ by a factor of 1.50 (2.21) in the self-calibration (known SR) scenario.
@@
In summary, we find CSST clusters can provide a constraint on $M_{\nu}$ of $\sim 0.1$ eV by combining N and P. With perfect knowledge of the SR parameters, this constraint can be tightened to $\sim 0.03$ eV, enabling a $2\sigma$ detection of the minimum neutrino mass of 0.06 eV. Recently, Lin et al. [90] presented a forecast using mock data from the CSST photometric galaxy clustering and cosmic shear surveys, which yielded an upper limit for $M_{\nu}$ of approximately $0.23$ eV at a 68% C.L. By combining clusters and galaxies, it is expected that CSST will significantly strengthen the constraint on the total neutrino mass. Moreover, the development of new probes such as voids [91] and Minkowski functionals [92, 93] has opened up new avenues for constraining neutrino masses. By combining these novel probes with traditional methods, it is expected that in the near future it will be possible to accurately constrain neutrino masses and distinguish between different mass hierarchies.
@@ Acknowledgements.
We thank Yan Gong for useful conversations. This work is supported by the National Key R&D Program of China Grant No. 2022YFF0503404 and No. 2021YFC2203100, by the National Natural Science Foundation of China Grants No. 12173036 and 11773024, by the China Manned Space Project “Probing dark energy, modified gravity and cosmic structure formation by CSST multi-cosmological measurements” and Grant No. CMS-CSST-2021-B01, by the Fundamental Research Funds for Central Universities Grants No. WK3440000004 and WK3440000005, by Cyrus Chun Ying Tang Foundations, and by the 111 Project for ”Observational and Theoretical Research on Dark Matter and Dark Energy” (B23042).